OpenStudy (anonymous):

Find the vertex, focus, directrix, and focal width of the parabola. x = 3y2

OpenStudy (anonymous):

The regular form for this problem would be: \[(y-k)^2 = 4a(x-h) \] To get the problem as close to that form as possible you would want to divide both sides by 3 so that you have: \[y^2 = \frac{ 1 }{ 3 } x\] The vertex would be (h,k). The focus is (h+a,k) The directrix is x=h-a The focal is the 4a part. To get a for the focus and the directrix you will divide the 1/3 by 4. Let me know if you need further help. :)